GlueX PID Meeting, June 11, 2010

From GlueXWiki
Jump to: navigation, search

Location, Time

Friday, June 11, 2010
2:00 pm EDT
JLab: CEBAF Center, Room F326/7
ESNet: 8542553
EVO: direct meeting link (good through 8/20)
Conference Group: 800-377-8846, participant code: 77438230
Slides: talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2010-06 on the JLab CUE (you have to be a member of the "halld" group). This directory is accessible from the web at .


  1. Announcements
  2. Minutes from the last meeting
  3. Time-of-Flight
    1. Prototype progress
    2. Status of the MOU
  4. Start Counter
    1. prototype progress
      • attenuation length measurements
  5. Target
    1. Status of the MOU
    2. Status of refrigerator plan
  6. Action Items


FSU: Paul Eugenio, Sasha Ostrovidov Jlab: Tom Carstens, Eugene Chudakov, Mark Ito (chair), Chris Keith, Sascha Somov, Simon Taylor, Tim Whitlatch, Beni Zihlmann

Review of minutes from the last meeting

We went over the minutes.

PMT purchase for TOF prototypes

Paul sent the quote, Mark submitted the purchase requisition, and now the purchase order is all set to go out, we just need a shipping address at FSU to attach to the PO.

Refrigerator for the target

We started from the entry in the minutes about the lab-wide review of helium demand.

  • Chris commented that if we do get a pulse tube refrigerator for cryotargets, that it actually might be able to service a frozen spin target. We did note, again, that such a target is not part of any current plan for Hall D.
  • Tim is in the middle of a cryogenic requirements review for the Hall and a decision on purchase of the recommended refrigerator is pending on completion of his work.


We spent a lot of time updating issues discussed at the last meeting.


  • Beni told us the LED system planned for the calorimeters cannot easily change the intensity of the light output, as would be convenient for time-walk calibration.
  • We need to talk to Elton Smith or Dan Carman about experience with the CLAS TOF nitorgen laser calibration system.
  • Right now, Paul is working on a LED pulser system aimed at developing experience with the prototype counters.
  • Stability of pmts is an issue, but not as critical as for the calorimeters.
  • Paul informed us that the time-walk calibration was repeated on a time scale of order a year.
  • Eugene told of us of a laser system that was being used in Hall A that may be available.
  • Tom reminded us that the CLAS electromagnetic calorimeter also has a nitrogen laser calibration scheme. Kevin Giovanetti developed that system.

Change of Thickness of Counters

  • Paul is exploring where the effective edge of the photocathode are with the Hamamatsu tubes.
    • PMT spec. states that the minimum cathode diameter is 46 mm
    • cutting an ERT disk to that diameter to obscure photocathode
    • plans is to look for residual signal
  • We talked about the light guide to PMT coupling scheme.
    • current plan: silicon cookies
    • initially glued used (still being used on one of the current counters)
    • Tim has done lay-outs that incorporate compression of the cookie
    • Eugene mentioned RTV glue as a possibility; joint can easily be broken with a razor blade


  • large sample of ESR DF2000 MA in hand at FSU
  • studies about to start
  • final light tight overwrapping: still waiting for delivery

Other TOF issues

Changing the Data Acquisition

  • Sasha has received TI board
  • Will be running with LED system to explore time-walk; naturally will result in trigger rate increase over cosmics; the TI will help here to correctly correlate ADC and TDC hits
  • Note: using a VME64 bridge (not using a CODA ROC)

New Counters

FSU is changing change the trigger to use a longer counter; this will improve the rate for cosmic triggers.


Progress will resume now that DOE grant work is done.

New Action Items

  1. Ask Elton about experience with the CLAS TOF laser calibration system.